# Non-commutative rings with infinitely many maximal subrings

### v1

@@ Abstract.
We study rings with infinitely (only finitely) many maximal subrings. We prove that if $M$ is a maximal left/right ideal of a ring $T$ which is not an ideal of $T$, and $R$ is the idealizer of $M$, then $T$ has at least $|R/M|+1$ maximal left/right ideals which are not an ideal of $T$; in particular $T$ has at least $|R/M|+1$ distinct maximal subrings. Moreover, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subrings or $T$ is a quasi duo ring with certain algebraic properties similar to commutative rings. We prove that for a simple ring $R$, the ring $R\times R$ has only finitely many maximal subrings if and only if $R$ is finite. Also we study rings which are integral over their centers and have only finitely many maximal subrings. We prove that if $T$ is integral over its center and $T$ has more than $2^{\aleph_{0}}$ maximal (left/right) ideals, then $T$ has infinitely many maximal subrings. In particular, we see that if a $J$-semisimple ring $T$ is integral over its center and has only finitely many maximal subrings, then $T$ embeds in $S\times\prod_{i\in I}E_{i}$, where each $E_{i}$ is an absolutely algebraic field and $S$ is a finite semisimple ring. We see that if $T$ is a left Noetherian algebraic $K$-algebra over an infinite field $K$ and $T$ has only finitely many maximal subrings, then $T$ is a countable left Artinian ring which is integral over $\mathbb{Z}_{p}$, where $p=Char(K)$. We exactly determine when $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings. We see that if $R$ is an infinite Artinian ring, then $\mathbb{M}_{n}(R)$, $n>1$, and $R\times R$ have infinitely many maximal subrings.
@@
maximal subring, non-commutative ring, idealizer, left maximal ideal, integrality, direct product
@@ 1. Introduction
Let $T$ be an associative ring with identity $1\neq 0$ and $RgMax(T)$ be the set of all maximal subrings (which contain $1_{T}$) of $T$, then we study when $RgMax(T)$ is finite/infinite. In [2], the author characterized fields and affine rings which have only finitely many maximal subrings; and in [10], the authors studied commutative rings with infinitely/only finitely many maximal subrings. We present some facts from them (which are needed in this paper) as follows:
@@
    A field $E$ has only finitely many maximal subrings if and only if $E$ has a subfield $F$ without maximal subrings with $[E:F]$ is finite; if and only if every descending chain $\cdots\subset R_{2}\subset R_{1}\subset R_{0}=E$, where each $R_{i}$ is a maximal subring of $R_{i-1}$, $i\geq 1$, is finite; moreover $F$ is unique, all such chains have a same length, say $m$, and $R_{m}=F$, see [2, Theorems 3.2 and 3.6].
@@
    Let $R=F[\alpha_{1},\ldots,\alpha_{n}]$ be an affine integral domain over a field $F$. Then $R$ has only finitely many maximal subrings if and only if $F$ has only finitely many maximal subrings and each $\alpha_{i}$ is algebraic over $F$, see [2, Theorem 4.1].
@@
    For each ring $R$, either $R$ is a Hilbert ring or $RgMax(R)$ is infinite, see [10, Corollary 1.9].
@@
    If $R$ is an Artinian ring which is either uncountable or of zero characteristic and or is not integral over its prime subring, then $RgMax(R)$ is infinite, see [10, Theorem2.5].
@@
    If $R$ is a Noetherian ring with $|R|>2^{\aleph_{0}}$, then $|RgMax(R)|\geq 2^{\aleph_{0}}$, see [10, Theorem 2.7].
@@
    If $\{R_{i}\ |\ i\in I\}$ is a family of rings and $R=\prod_{i\in I}R_{i}$, then in [10, Theorem 3.9], the authors exactly determined when $R$ has only finitely many maximal subrings. Moreover, if $I$ is infinite then $|RgMax(R)|\geq 2^{|I|}$, see [9, Remark 3.18].
@@
    If $K$ is a field, then $K\times K$ has finitely many maximal subrings if and only if $K$ is a finite field, see [10, Corollary 3.5].
@@
    If $R$ is a semilocal ring, then either $R$ has infinitely many maximal subrings or $R$ is integral over $\mathbb{Z}_{n}$, for some $n>1$, see [10, Corollary 2.2].
@@
    If $R$ is a ring with $|Max(R)|>2^{\aleph_{0}}$ (resp. $|R/J(R)|>2^{2^{\aleph_{0}}}$), then $|RgMax(R)|\geq|Max(R)|$ (resp. $|RgMax(R)|\geq 2^{\aleph_{0}}$), see [10, Proposition2.8].
@@
    If $R$ is a ring, then $R\times R$ has only finitely many maximal subrings if and only if $R$ is a semilocal ring with only finitely many maximal subrings and $R/M$ is finite for each $M\in Max(R)$, see [10, Corollary 3.10].
@@
Finally in [10], the authors generalized $(1)$ for a finite direct products of fields too. For non-commutative rings, the existence of maximal subring first studied in [1], we will mentioned some of the result of [1] in the next section. It is interesting to know that each ring $R$ can be considered as a maximal subring of a larger ring $T$, see [1, Theorem 3.7]. Recently, in [5] and [6] the existence and the structure of maximal subrings in division rings and the existence of maximal subrings in certain non-commutative rings investigated, respectively. In [20] and [7], the authors systematically studied maximal commutative subrings for general noncommutative rings and commutative maximal subrings, respectively. We refer the interested reader to [24], for classification of maximal subrings of finite semisimple rings and their applications to the calculation of the covering number of a finite semisimple ring; and also to [3, 4] for more results about the conductor ideals and certain ideals of the extension $R\subseteq T$, where $R$ is a maximal subring of $T$.
@@
A brief outline of this paper is as follows. In Section $2$, we give some preliminaries results that needed in this paper, specially about the number of maximal left/right ideals of a ring $T$ which are not ideals of $T$. In fact, we prove that if $T$ is a ring and $M$ is a maximal left/right ideal of $T$ which is not an ideal of $T$ and $R$ is the idealizer of $M$ in $T$, then $T$ has at least $|R/M|+1$ maximal left/right ideals $N$ such that $N$ is similar to $M$ (i.e., $T/N\cong T/M$ as left/right $T$-module) and therefore $T$ has at least $|R/M|+1$ maximal left/right ideals which are not ideals of $T$. Moreover, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ is a quasi duo ring or $T$ has infinitely many maximal left/right ideals which are not ideals of $T$ (in fact, either $|Max_{l}(T)\setminus Max(T)|\geq|K|$ or $|Max_{r}(T)\setminus Max(T)|\geq|K|$). Finally we present some needed results from [27] about the (number of) maximal left/right ideals of $\mathbb{M}_{n}(R)$, where $R$ is a ring and $n\in\mathbb{N}$, and also from [1] about maximal subrings of non-commutative rings.
@@
In Section $3$, we investigate on the number of maximal subrings in general non-commutative rings. First we see that if $D$ is a division ring with finitely many maximal subrings, then the center of $D$ is algebraically closed in $D$. For a simple ring $S$, we exactly determine all maximal subrings of $S\times S$. We see that if $S$ is a simple ring which is not a division ring, then $S$ has only finitely many maximal subrings if and only if $S$ is a finite ring. Consequently, if $S$ is a simple then $S\times S$ has only finitely many maximal subrings if and only if $S$ is a finite ring. Next, we see that if a ring $T$ has only finitely many maximal subrings, then $Max_{l}(T)\setminus Max(T)$, $Max_{r}(T)\setminus Max(T)$, $Max(T)\setminus Max_{l}(T)$, $Prm_{l}(T)\setminus Max(T)$ and $Prm_{l}(T)\setminus Max_{l}(T)$ (and analogously for right sets) are all finite, moreover if $M$ is a maximal left/right ideal of $T$ which is not an ideal of $T$, then $\mathbb{I}(M)/M$ is a finite field and if $N$ is a maximal ideal of $T$ which is not a maximal left ideal of $T$, then $T/N\cong\mathbb{M}_{n}(F)$, for some finite field $F$ and $n>1$. Also we see that if $T$ is a ring with only finitely many maximal subrings, then $J(T)=\bigcap_{M\in Max(T)}M$. In particular, if we assume that $T$ is a $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subrings or $T$ is a quasi duo ring. Finally in Section $3$, we prove that if $T$ is an infinite ring with nonzero characteristic, then either $T$ has infinitely many maximal subrings or $\{M\in Max(T)\ |\ |T/M|<\aleph_{0}\}$ is a finite set.
@@
In Section $4$, we studied rings which are integral over their centers and have only finitely many maximal subrings. In particular, algebraic $K$-algebra over an infinite field $K$. We prove that if $T$ is a ring which is integral over its center and $T$ has more than $2^{\aleph_{0}}$ maximal (left/right) ideals, then $T$ has infinitely many maximal subrings. We observe that, if $T$ is an algebraic $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subring or $T/J(T)$ is a commutative von Neumann regular ring which is integral over $\mathbb{Z}_{p}$, where $p=Char(K)$. We see that if $T$ is integral over its center, then either $T$ has infinitely many maximal subrings or $Prm_{l}(T)=Max(T)=Prm_{r}(T)$, and for each maximal ideal $M$ of $T$, either $T/M$ is an absolutely algebraic field (with only finitely many maximal subrings) or $T/M$ is a finite simple ring; in particular, if in addition $J(T)=0$, then $T$ embeds in $S\times\prod_{i\in I}E_{i}$, where $S$ is a finite semisimple ring and each $E_{i}$ is an absolutely algebraic field.
@@
Finally in Section $5$, we exactly determine when $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings. We see that if $R$ is an infinite Artinian ring, then $\mathbb{M}_{n}(R)$, $n>1$, and $R\times R$ have infinitely many maximal subrings.
@@
All rings in this paper are unital with $1\neq 0$. All subrings, modules and homomorphisms are also unital. If $R\subsetneq T$ is a ring extension and there exists no other subring between $R$ and $T$, then $R$ is called a maximal subring of $T$, or the extension $R\subseteq T$ is called a minimal ring extension, see [14, 15, 16, 18, 25]. If $T$ is a ring, $I$ is an ideal of $T$ and $M$ is a left (resp. right) $T$-module, then $Min_{T}(I)$, $Max_{r}(T)$, $Max_{l}(T)$, $Max(T)$ and $l.ann_{T}(M)$ (resp. $r.ann_{T}(M)$), denote the set of all minimal prime ideals of $I$ in $T$, the set of all maximal right ideals of $T$, the set of all maximal left ideals of $T$, the set of all maximal ideals of $T$ and the left annihilator of $M$ in $T$ (resp. the right annihilator of $M$ in $T$), respectively. We use $Min(T)$ for $Min_{T}(0)$. The set of all left (resp. right) primitive ideals of $T$ is denoted by $Prm_{l}(T)$ (resp. $Prm_{r}(T)$). The characteristic of a ring $T$ is denoted by $Char(T)$. If $T$ is a ring, then $dim(T)$ denotes the classical Krull dimension of $T$ (i.e., the supremum of the lengths of all chains of prime ideals of $T$). Also we use $Aut(T)$ for the set of all ring automorphism of a ring $T$. $v.dim_{K}(T)$ denotes the vector dimension of a $K$-algebra $T$ over a field $K$. $J(T)$ denotes the Jacobson radical of a ring $T$. For a ring $T$, $Nil_{*}(T)$, $Nil^{*}(T)$, $N(T)$ and $U(T)$ are the lower nilradical, the upper nilradical, the set of all nilpotent elements and the units group of $T$, respectively. If $T$ is a ring, the opposite ring of $T$ is denoted by $T^{op}$. A prime ideal $P$ of a ring $T$ is called strongly (resp. completely) prime if $T/P$ has no nonzero nil ideal (resp. $T/P$ is a domain). If $T$ is a ring, then $\mathbb{M}_{n}(T)$ denote the ring of all $n\times n$ square matrices over $T$. If $M$ is a left (resp. right) module over a ring $T$, then the ring of all $T$-module homomorphisms of $M$ is denoted by $End({}_{T}M)$ (resp. $End(M_{T})$). If $A$ is a right ideal of a ring $T$, then $\mathbb{I}(A)$ denotes the idealizer of $A$ in $T$, that is the largest subring of $T$ which $A$ is a two-sided ideal in it, i.e., $\mathbb{I}(A)=\{t\in T\ |\ tA\subseteq A\}$, the similar notation is used when $A$ is a left ideal of $T$. If $A$ is a right ideal of a ring $T$, then $\mathbb{E}_{T}(A):=\mathbb{I}(A)$ is called the eigenring of $A$. If $X$ is a subset of a ring $T$, then $C_{T}(X)$ is the centralizer of $X$ in $T$, in particular, $C(T)=C_{T}(T)$ is the center of $T$. For a ring $T$, the subring $\{n.1_{T}\ |\ n\in\mathbb{Z}\}$ is called the prime subring of $T$. It is clear that $T$ and $T^{op}$ have a same prime subring (moreover, $C(T)=C(T^{op})$) and therefore have a same characteristic. A ring $T$ is called left (resp. right) quasi duo if each maximal left (resp. right) ideal of $T$ is a two-sided ideal of $T$, see [22]. $T$ is called quasi duo, if $T$ is left and right quasi duo ring. A ring $T$ is called left (resp. right) duo if each left (resp. right) ideal of $T$ is two-sided. Similarly duo rings are defined. If $R$ is a subring of a ring $T$ and $t\in T$, then we say that $t$ is left (resp. right) integral over $R$, if $t$ is a root of a left (resp. right) monic polynomial of degree $n\geq 1$ over $R$. $R$ is called left (resp. right) integrally closed in $T$, if every left (resp. right) integral element of $T$ over $R$, belongs to $R$. $R$ is called integrally closed in $T$, whenever $R$ is left and right integrally closed in $T$. If $T$ is a ring and $R$ is a subring of the center of $T$, then $T$ is called an $R$-algebra. It is clear that, if $T$ is an $R$-algebra, then the definitions of right and left integral elements and also left or right integrally closed for the extension $R\subseteq T$ are coincided. If $F\subseteq E$ is a field extension, then $tr.deg(E/F)$ denotes the transcendence degree of $E$ over $F$. For other notations and definitions we refer the reader to [19, 21, 26].

### v2

# Noncommutative rings with infinitely many maximal subrings

@@ Abstract.
We investigate rings with infinitely many or only finitely many maximal subrings. We prove that if $M$ is a maximal left/right ideal of a ring $T$ that is not an ideal of $T$, and $R$ is the idealizer of $M$, then $T$ has at least $|R/M|+1$ maximal left/right ideals that are not two-ideals of $T$; in particular $T$ has at least $|R/M|+1$ distinct maximal subrings. Moreover, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subrings or $T$ is a quasi-duo ring with certain algebraic properties similar to those commutative rings that have not infinitely many maximal subrings. We prove that for a simple ring $R$, the ring $R\times R$ has only finitely many maximal subrings if and only if $R$ is finite. We also study rings which are integral over their centers and have only finitely many maximal subrings. We prove that if $T$ is integral over its center and $T$ has more than $2^{\aleph_{0}}$ maximal (left/right) ideals, then $T$ has infinitely many maximal subrings. In particular, we show that if a $J$-semisimple ring $T$ is integral over its center and has only finitely many maximal subrings, then $T$ embeds into $S\times\prod_{i\in I}E_{i}$, where each $E_{i}$ is an absolutely algebraic field and $S$ is a finite semisimple ring. Furthermore, if $T$ is a left Noetherian algebraic $K$-algebra over an infinite field $K$ and $T$ has only finitely many maximal subrings, then $T$ is a countable left Artinian ring which is integral over $\mathbb{Z}_{p}$, where $p=\operatorname{Char}(K)$. We precisely determine when $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$, where each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$, has only finitely many maximal subrings. Finally, we see that if $R$ is an infinite Artinian ring, then $\mathbb{M}_{n}(R)$, for $n>1$, and $R\times R$ have infinitely many maximal subrings.
@@
maximal subring, noncommutative ring, idealizer, maximal left ideal, integrality, direct product
@@ 1. Introduction
Let $T$ be an associative ring with identity $1\neq 0$ and let $\operatorname{RgMax}(T)$ denote the set of all maximal subrings (which contain $1_{T}$) of $T$. In this paper we study conditions under which $\operatorname{RgMax}(T)$ is finite or infinite. In [2], the author characterized fields and affine rings that have only finitely many maximal subrings; in [10], the authors studied commutative rings with either infinitely many or only finitely many maximal subrings. We recall some relevant facts from those works (which will be needed later) as follows:
@@
    A field $E$ has only finitely many maximal subrings if and only if it contains a subfield $F$ without maximal subrings with $[E:F]$ is finite; equivalently, every descending chain $\cdots\subset R_{2}\subset R_{1}\subset R_{0}=E$, where each $R_{i}$ is a maximal subring of $R_{i-1}$, $i\geq 1$, is finite. Moreover, such a subfield $F$ is unique, all such chains have the same length, say $m$, and $R_{m}=F$; see [2, Theorems 3.2 and 3.6].
@@
    Let $R=F[\alpha_{1},\ldots,\alpha_{n}]$ be an affine integral domain over a field $F$. Then $R$ has only finitely many maximal subrings if and only if $F$ has only finitely many maximal subrings and each $\alpha_{i}$ is algebraic over $F$; see [2, Theorem 4.1].
@@
    For any ring $R$, either $R$ is a Hilbert ring or $\operatorname{RgMax}(R)$ is infinite, see [10, Corollary 1.9].
@@
    If $R$ is an Artinian ring that is either uncountable, or of zero characteristic, or is not integral over its prime subring, then $\operatorname{RgMax}(R)$ is infinite; see [10, Theorem2.5].
@@
    If $R$ is a Noetherian ring with $|R|>2^{\aleph_{0}}$, then $|\operatorname{RgMax}(R)|\geq 2^{\aleph_{0}}$; see [10, Theorem 2.7].
@@
    Let $\{R_{i}\ |\ i\in I\}$ be a family of rings and $R=\prod_{i\in I}R_{i}$, in [10, Theorem 3.9], the authors precisely determined when $R$ has only finitely many maximal subrings. Moreover, if $I$ is infinite, then $|\operatorname{RgMax}(R)|\geq 2^{|I|}$; see [9, Remark 3.18].
@@
    For a field $K$, the ring $K\times K$ has only finitely many maximal subrings if and only if $K$ is a finite field; see [10, Corollary 3.5].
@@
    If $R$ is a semilocal ring, then either $R$ has infinitely many maximal subrings or $R$ is integral over $\mathbb{Z}_{n}$, where $\operatorname{Char}(R)=n>1$; see [10, Corollary 2.2].
@@
    If $R$ is a ring with $|\operatorname{Max}(R)|>2^{\aleph_{0}}$ (resp. $|R/J(R)|>2^{2^{\aleph_{0}}}$), then $|\operatorname{RgMax}(R)|\geq|\operatorname{Max}(R)|$ (resp. $|\operatorname{RgMax}(R)|\geq 2^{\aleph_{0}}$); see [10, Proposition2.8].
@@
    A ring $R$ satisfies that $R\times R$ has only finitely many maximal subrings if and only if $R$ is a semilocal ring with only finitely many maximal subrings and $R/M$ is finite for each $M\in\operatorname{Max}(R)$; see [10, Corollary 3.10].
@@
In [10], the authors also generalized statement $(1)$ for a finite direct products of fields. For noncommutative rings, the existence of maximal subring was first studied in [1]; we will mention some results from [1] in the next section. Interestingly, every ring $R$ can be realized as a maximal subring of some larger ring $T$; see [1, Theorem 3.7]. Recently, the existence and the structure of maximal subrings in division rings were investigated in[5]; and the existence of maximal subrings in certain noncommutative rings was studied in [6]. In [20] and [7], the authors systematically examined maximal commutative subrings of general noncommutative rings and commutative maximal subrings, respectively. We refer the interested reader to [24] for classification of maximal subrings of finite semisimple rings and their applications to calculating the covering number of a finite semisimple ring; see also [3, 4] for further results on conductor ideals and certain ideals of the extension $R\subseteq T$, where $R$ is a maximal subring of $T$.
In this paper, we extend the study of maximal subring beyond the commutative setting, focusing on general associative rings with identity. Specifically, we investigate conditions under which a ring possesses only finitely many or infinitely many maximal subrings, with particular attention to noncommutative algebras over infinite fields, rings integral over their centers, certain matrix or product constructions. Our approach combines ideal-theoretic techniques, structural ring theory, and cardinality arguments, providing a unified framework that generalizes and complements earlier commutative results.
@@
A brief outline of this paper is as follows. In Section $2$, we present some preliminary results needed in the sequel, particularly concerning the number of maximal left/right ideals of a ring $T$ that are not two-sided ideals. Specifically, we prove that if $T$ is a ring, $M$ is a maximal left/right ideal of $T$ that is not an ideal of $T$, and $R$ is the idealizer of $M$ in $T$, then $T$ has at least $|R/M|+1$ maximal left/right ideals $N$ that are similar to $M$ (i.e., $T/N\cong T/M$ as left/right $T$-modules); consequently, $T$ has at least $|R/M|+1$ maximal left/right ideals that are not ideals of $T$. Moreover, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ is a quasi-duo ring or $T$ has infinitely many maximal left/right ideals that are not ideals of $T$ (indeed, either $|\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)|\geq|K|$ or $|\operatorname{{Max}_{r}}(T)\setminus\operatorname{Max}(T)|\geq|K|$). Finally we recall several required results from [27] regarding the number of maximal left/right ideals of $\mathbb{M}_{n}(R)$ (where $R$ is a ring and $n\in\mathbb{N}$), as well as from [1] concerning maximal subrings of noncommutative rings.
@@
In Section $3$, we investigate the number of maximal subrings in general noncommutative rings. First, we show that if $D$ is a division ring with only finitely many maximal subrings, then the center of $D$ is algebraically closed in $D$. For a simple ring $S$, we precisely determine all maximal subrings of $S\times S$. We prove that if $S$ is a simple ring that is not a division ring, then $S$ has only finitely many maximal subrings if and only if $S$ is finite. Consequently, for a simple ring $S$, the ring $S\times S$ has only finitely many maximal subrings if and only if $S$ is a finite ring. Next, we demonstrate that if a ring $T$ has only finitely many maximal subrings, then $\operatorname{{Max}_{\ell}}(T)\setminus\operatorname{Max}(T)$, $\operatorname{{Max}_{r}}(T)\setminus\operatorname{Max}(T)$, $\operatorname{Max}(T)\setminus\operatorname{{Max}_{\ell}}(T)$, $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{Max}(T)$ and $\operatorname{{Prm}_{\ell}}(T)\setminus\operatorname{{Max}_{\ell}}(T)$ (and their right analogous) are all finite. Furthermore, if $M$ is a maximal left/right ideal of $T$ that is not an ideal of $T$, then $\mathbb{I}(M)/M$ is a finite field; and if $N$ is a maximal ideal of $T$ that is not a maximal left ideal of $T$, then $T/N\cong\mathbb{M}_{n}(F)$, for some finite field $F$ and some $n>1$. We also show that, if $T$ has only finitely many maximal subrings, then $J(T)=\bigcap_{M\in\operatorname{Max}(T)}M$. In particular, if $T$ is a $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subrings or $T$ is a quasi-duo ring. Finally in Section $3$, we prove that if $T$ is an infinite ring of nonzero characteristic, then either $T$ has infinitely many maximal subrings or the set $\{M\in\operatorname{Max}(T)\ |\ |T/M|<\aleph_{0}\}$ is finite.
@@
Section $4$, is devoted to rings that are integral over their centers and have only finitely many maximal subrings, with special attention to algebraic $K$-algebra over an infinite field $K$. We prove that if $T$ is integral over its center and $T$ has more than $2^{\aleph_{0}}$ maximal (left/right) ideals, then $T$ has infinitely many maximal subrings. Moreover, if $T$ is an algebraic $K$-algebra over an infinite field $K$, then either $T$ has infinitely many maximal subring or $T/J(T)$ is a commutative von Neumann regular ring that is integral over $\mathbb{Z}_{p}$, where $p=\operatorname{Char}(K)$. We also establish that if $T$ is integral over its center, then either $T$ has infinitely many maximal subrings or $\operatorname{{Prm}_{\ell}}(T)=\operatorname{Max}(T)=\operatorname{{Prm}_{r}}(T)$, and for each maximal ideal $M$ of $T$, either $T/M$ is an absolutely algebraic field (with only finitely many maximal subrings) or $T/M$ is a finite simple ring. In particular, if additionally $J(T)=0$, then $T$ embeds into $S\times\prod_{i\in I}E_{i}$, where $S$ is a finite semisimple ring and each $E_{i}$ is an absolutely algebraic field.
@@
Finally in Section $5$, we precisely determine when a ring of the form $T=\prod_{i\in I}\mathbb{M}_{n_{i}}(E_{i})$ (with each $E_{i}$ is a field and $n_{i}\in\mathbb{N}$), has only finitely many maximal subrings. We also prove that if $R$ is an infinite Artinian ring, then $\mathbb{M}_{n}(R)$, for $n>1$, as well as $R\times R$, have infinitely many maximal subrings.
@@
All rings considered are unital with $1\neq 0$. All subrings, modules and homomorphisms are assumed to be unital. If $R\subsetneq T$ is a ring extension and such that no other subring lies strictly between $R$ and $T$, then $R$ is called a maximal subring of $T$; equivalently, the extension $R\subseteq T$ is called a minimal ring extension, see [14, 15, 16, 18, 25]. Let $T$ be a ring and $M$ be a left $T$-module, then $\operatorname{{Max}_{r}}(T)$, $\operatorname{{Max}_{\ell}}(T)$, $\operatorname{Max}(T)$ and $\operatorname{l.ann}_{T}(M)$, denote the set of all maximal right ideals of $T$, the set of all maximal left ideals of $T$, the set of all maximal ideals of $T$ and the left annihilator of $M$ in $T$, respectively. The set of all left (resp. right) primitive ideals of $T$ is denoted by $\operatorname{{Prm}_{\ell}}(T)$ (resp. $\operatorname{{Prm}_{r}}(T)$). The characteristic of a ring $T$ written as $\operatorname{Char}(T)$. We use $\operatorname{Aut}(T)$ for the group of all ring automorphisms of $T$. The Jacobson radical of $T$ is denoted by $J(T)$. We let $N(T)$ and $U(T)$ be the set of all nilpotent elements and the units group of $T$, respectively. The opposite ring of $T$ is written by $T^{\text{op}}$. The ring of all $n\times n$ matrices over $T$ is denoted by $\mathbb{M}_{n}(T)$. The ring of all $T$-module homomorphisms of $M$ is written as $\operatorname{End}_{T}(M)$. For a right ideal $A$ of $T$, the idealizer of $A$ in $T$ is denoted by $\mathbb{I}(A)$, it is the largest subring of $T$ in which $A$ is a two-sided ideal, i.e., $\mathbb{I}(A)=\{t\in T\ |\ tA\subseteq A\}$. The same notation is used when $A$ is a left ideal of $T$. In any case, the ring $\mathbb{E}_{T}(A):=\mathbb{I}(A)/A$ is called the eigenring of $A$. If $X$ is a subset of $T$, then $C_{T}(X)$ denotes the centralizer of $X$ in $T$; in particular, $C(T)=C_{T}(T)$ is the center of $T$. The subring $\{n.1_{T}\ |\ n\in\mathbb{Z}\}$ is called the prime subring of $T$. A ring $T$ is called left (resp. right) quasi-duo if any maximal left (resp. right) ideal of $T$ is two-sided. If $T$ is both left and right quasi-duo, it is called a quasi-duo ring, see [22]. Let $R$ be a subring of $T$ and $t\in T$. We say that $t$ is left (resp. right) integral over $R$, if $t$ is a root of a monic polynomial of degree $n\geq 1$ with coefficients on the left (resp. right) side from $R$. The subring $R$ is left (resp. right) integrally closed in $T$, if every element of $T$ that is left (resp. right) integral element over $R$, belongs to $R$. We say it is integrally closed in $T$, if $R$ is both left and right integrally closed in $T$. When $T$ is a ring and $R$ is a subring of $C(T)$, we call $T$ is an $R$-algebra. In this case, the notation of right and left integral elements (as well as, left and right integrally closed) for the extension $R\subseteq T$ coincide. For a field extension $F\subseteq E$, we write $\operatorname{tr.deg}(E/F)$ for the transcendence degree of $E$ over $F$. Other notations and definitions follow [19, 21, 26].